OpenStudy (anonymous):

Find x in the following right triangle, where a = 9, and b = 12.

OpenStudy (anonymous):

Do I need to post the picture too? @hartnn

hartnn (hartnn):

yes

OpenStudy (anonymous):

Picture:

OpenStudy (anonymous):

21-180=159 degrese

hartnn (hartnn):

use pythagoras theorem

hartnn (hartnn):

 \(\huge x^2=a^2+b^2\)

hartnn (hartnn):

\(x^2=9^2+12^2\)

hartnn (hartnn):

find x

OpenStudy (anonymous):

225?

hartnn (hartnn):

thats x^2 whats x ?

OpenStudy (anonymous):

How do I find x?

hartnn (hartnn):

\(x^2= 225 \\ x = \sqrt {225}\)

OpenStudy (anonymous):

oooooooooh ok

hartnn (hartnn):

use calculator if you already don't know

OpenStudy (anonymous):

15

hartnn (hartnn):

correct! :)

OpenStudy (anonymous):

I always use a calculator lol

OpenStudy (anonymous):

thanks !